# Counterfactual Multi-player Bandits for Explainable Recommendation Diversification

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systemsccs: Computing methodologies Multi-agent reinforcement learning
@@ 1. Introduction
Recommendation systems (RS) are widely deployed in various online platforms, such as Google, Facebook, and Yahoo!, to mitigate information overload. However, most existing recommendation methods (Rendle et al., 2009; He et al., 2020; Koren et al., 2009; Cheng et al., 2016) only prioritize recommending the most relevant items to users, which can have negative consequences for both users and service providers. Users may experience the “filter bubble” (Pariser, 2011) problem, receiving recommendations based solely on their past behavior, leading to a lack of content diversity and reduced long-term satisfaction. Content providers may face the “Matthew Effect” (Merton, 1968), where new or niche content/products yield less exposure, discouraging the engagement of content providers. Therefore, it is crucial to consider not only relevance but also recommendation diversity to enhance the overall user experience and maintain a healthy ecosystem for content providers.
@@ 2.1. Recommendation Diversification
To address the “filter bubble” and reduced provider engagement issues, it is crucial to recommend accurate and diverse items for a healthier online marketplace. Existing diversification methods are typically offline and categorized as pre-processing, in-processing, and post-processing methods (Wu et al., 2024). Pre-processing methods (Kwon et al., 2020; Cheng et al., 2017; Paudel et al., 2017; Zheng et al., 2021) involve preparing interaction data before model training. In-processing integrates diversity into the training process, using it as regularization (Chen et al., 2020; Maropaki et al., 2020; Sanz-Cruzado and Castells, 2018; Wasilewski and Hurley, 2016; Zhou et al., 2020) or a ranking score (Li et al., 2017; Yu, 2022). Post-processing, the most scalable, includes greedy-based methods like MMR (Carbonell and Goldstein, 1998; Santos et al., 2010; Vargas et al., 2011) and DPP (Gan et al., 2020; Chen et al., 2018; Gong et al., 2022; Huang et al., 2021; Wilhelm et al., 2018), which adjust item selection and rankings to balance relevance and diversity, and refinement-based methods (Li et al., 2020b; Tsukuda and Goto, 2019), which modify positions or replace items based on diversity metrics. Other online methods, such as bandit strategies (Ding et al., 2021; Li et al., 2020a; Qin et al., 2014; Parapar and Radlinski, 2021), treat diversity as part of the score on each arm (item or topic) in the bandit recommendation algorithms and reinforcement learning (Stamenkovic et al., 2022; Zheng et al., 2018), continuously update based on user feedback for long-term optimization.
@@ 2.2. Explainable Recommendation
Explainable recommendations have attracted significant attention in academia and industry, aiming to enhance transparency, user satisfaction, and trust (Zhang et al., 2020, 2014). Early methods focused on generating individualized explanations, often customizing models and using auxiliary information (Zhang et al., 2014; Lu et al., 2018; Wu et al., 2019). For example, the Explicit Factor Model (EFM) (Zhang et al., 2014) recommends products based on features extracted from user reviews. Other approaches decouple explanations from the recommendation model, making them post-hoc and model-agnostic (Ribeiro et al., 2016; Wang et al., 2018; Singh and Anand, 2019). Recently, counterfactual reasoning has been widely used to improve explainability (Wachter et al., 2017; Ghazimatin et al., 2020; Tan et al., 2021; Xiong et al., 2021; Ge et al., 2022). For instance, PRINCE (Ghazimatin et al., 2020) identifies minimal user actions that, if removed, would change the recommendation, while CEF (Ge et al., 2022) uses counterfactual reasoning to explain fairness in feature-aware recommendation systems.
@@ 3.2. Counterfactual Framework for Explainable Diversification
Figure 1. The architecture of CMB. CMB consists of three major stages: the first stage of base model training, the second stage of counterfactual framework within multi-player bandits optimization, and the third stage of explanation.
@@ 3.3. Multi-player Bandits for Diversity Optimization
When all players have selected the arm $\Delta_{p}^{t}$ , we can get the $\Delta$ and further the reward $r_{t}$ (Eq. 9) based on counterfactual result ~$\widetilde{\mathcal{R}}_{g}$. To reduce the computational complexity, we update the cumulative reward value $V^{t+1}_{\Delta_{p}}$ of the corresponding arm $\Delta_{p}$ selected by each player $p$ in iteration $t+1$ by an incremental average method:
(13) $V^{t+1}_{\Delta_{p}}=V^{t}_{\Delta_{p}}+\frac{1}{n}[r_{t}-V^{t}_{\Delta_{p}}],$
@@
Discussion. (1) We mainly work on latent features of items to explain the recommendation diversification. The key consideration is that current mainstream recommendation models still originate from collaborative filtering methods, which are based on latent features. These motivate us to work on the latent features for controlling the diversity level of recommendation results. It is worth noting that our model can both work raw and latent features. (2) The main reason why we decided to apply the perturbation $\Delta$ to the features of items is that perturbing at the items’ feature level enables us to identify specific features that impact the diversity of the model at a more fine-grained feature level. (3) The multi-armed bandits method, unlike previous approaches, is primarily utilized in online recommendation scenarios. In our study, we employ this method to optimize the learning objective in the counterfactual framework, especially targeting and optimizing the non-differentiable diversity metrics directly.
@@ 4.1.1. Datasets
We performed experiments on three widely used real-world datasets, MovieLens 1M (Harper and Konstan, 2015) (ML1M), MovieLen 10M (Harper and Konstan, 2015) (ML10M), and Amazon CDs and Vinyl (Ni et al., 2019) (CDs) to evaluate the models under different data scales, data sparsity, and application scenarios. The statistics of the datasets are shown in Table 2.
@@ 4.1.4. Implementation Details
In our experiments, we employ BPRMF and LightGCN as the base model $g$, with LightGCN set to three layers of the graph neural network. For baselines, BPRMF calculates the relevance scores for MMR and constructs the kernel matrix for DPP. The trade-off parameter for MMR and DPP is empirically set to 0.9 after testing values from {0.1, 0.3, 0.5, 0.7, 0.9}, prioritizing guaranteeing the optimal accuracy performance as much as possible. All models have a latent feature dimension of 50, with $K$ set to 20 for counterfactual learning. The $\alpha$ in $\alpha$-nDCG is set to 0.5 same as (Santos et al., 2010; Vargas et al., 2011; Li et al., 2017; Qin et al., 2020; Yu, 2022), and model parameters are optimized by Adam (Kingma and Ba, 2014) with a learning rate of 0.005. In the bandit algorithm, we set the thrfeshold of arm values $A$ to 0.3, the number of arms $n_{A}$ to 61, and $\epsilon$ to 0.1. The hyperparameters $\lambda_{1}$, $\lambda_{2}$, and the total number of iterations $T$ are set to 5, 0.9, and 200, respectively. The source code is available at https://github.com/Forrest-Stone/CMB.
@@
Table 3. Comparisons of the accuracy and diversity performance. The bold scores are the best in each column, and the underlined scores are the second best. The symbols $\uparrow$ and $\downarrow$, along with their preceding values, represent the percentage (is omitted) improvement and decrease of a given method in the corresponding metric, in comparison to the base model $g$.
@@ 4.2. Performance Comparison
Second, not only the combined optimization objective helps achieve a reasonable balance between accuracy and diversity, but also the single diversity objective does. For instance, as shown in Table 5, when compared to BPRMF, $\text{CMB}_{\text{BPRMF}}$-ILAD shows a decrease of 18.79% in Recall@10 and an increase of 40.62% in ILAD@10 on ML10M, while showing a decrease of 31.29% in NDCG@10 and an increase of 59.58% in PC@10 on CDs. These results demonstrate that the proposed single diversity objective can improve diversity performance while maintaining accuracy performance, and outperforming other diversity methods like MMR and DPP.
@@
Table 4. Comparisons of the accuracy and diversity performance. The bold scores are the best in each column, and the underlined scores are the second best. The symbols $\uparrow$ and $\downarrow$, along with their preceding values, represent the percentage (is omitted) improvement and decrease of a given method in the corresponding metric, in comparison to the base model $g$.
@@
Table 5. Comparisons among the accuracy and diversity performance of CMB optimizes the single diversity metrics. $\text{CMB}_{\text{ILAD}}^{\text{Gradient}}$ represents CMB that directly optimizes the differentiable metric ILAD by using the gradient method. $\text{CMB}_{\text{BPRMF}}$-Random represents CMB that chooses the arm randomly for each player. The bold scores are the best in each column.
@@
Second, as illustrated in Tables 3 and 5, $\text{CMB}_{\text{BPRMF}}$-SC achieves the second-best SC@10 on ML10M, only 0.12% decrease compared to the best baseline. Similarly, $\text{CMB}_{\text{BPRMF}}$-PC also achieves the second-best performance in the PC metric on ML10M and CDs. These results demonstrate that CMB can achieve a satisfactory diversification. Moreover, while CMB sometimes trails DPP in diversity metrics, this is beacuse DPP prioritizes diversity over accuracy. In contrast, our method balances both, leading to improved accuracy even if diversity scores are slightly lower when compared to the DPP.
Application observations. First, as shown in Table 3 and 4, CMB can significantly improve various base models. Notably, when optimizing $\alpha$-nDCG, $\text{CMB}_{\text{BPRMF}}$ outperforms the base model itself, demonstrating the robustness of CMB. Due to its model-independent property, we can select a suitable base model $g$ to maximize the appropriate diversity metrics based on our specific demands, enabling us to handle a wide range of application scenarios.
Second, from Table 5, by comparing $\text{CMB}_{\text{BPRMF}}^{\text{Gradient}}$-ILAD and $\text{CMB}_{\text{BPRMF}}$-ILAD, we observe that the gradient descent optimization method outperforms the bandit optimization method in terms of accuracy on CDs, while exhibiting better diversity results on ML10M for PC and ILAD metrics. Thus, the choice of which optimization method should be based on the specific application scenario (e.g., dataset, diversity level, etc.).
Finally, as shown in Table 3 and 5, compared to $\text{CMB}_{\text{Random}}$, CMB that only optimizes a single diversity metric also achieves good results on that metric, which shows that our bandit method is effective for optimizing different diversity metrics. Moreover, optimizing the combination of trade-off objectives simultaneously also achieves a good balance in accuracy and diversity, further highlighting the effectiveness of our combined optimization objectives. For example, NDCG@10 of $\text{CMB}_{\text{BPRMF}}$-SC-NDCG on ML10M and CDs is 18.88% and 34.08% higher than $\text{CMB}_{\text{BPRMF}}$-Random, respectively, while the SC@10 only decreases by 1.35% and 6.16%.
@@
4.3. Validity Analysis of Explanations
@@
4.4. Case Study of Explanations
@@ A.2. Performance Comparison
Table 7. Comparisons of the accuracy and diversity performance. The bold scores are the best in each column, and the underlined scores are the second best. The symbols $\uparrow$ and $\downarrow$, along with their preceding values, represent the percentage (is omitted) improvement and decrease of a given method in the corresponding metric, in comparison to the base model $g$.
@@
Table 8. Comparisons of the accuracy and diversity performance. The bold scores are the best in each column, and the underlined scores are the second best. The symbols $\uparrow$ and $\downarrow$, along with their preceding values, represent the percentage (is omitted) improvement and decrease of a given method in the corresponding metric, in comparison to the base model $g$.
@@
Table 9. Comparisons among the accuracy and diversity performance of CMB optimizes the single diversity metrics. $\text{CMB}_{\text{ILAD}}^{\text{Gradient}}$ represents CMB that directly optimizes the differentiable metric ILAD by using the gradient method. $\text{CMB}_{\text{BPRMF}}$-Random represents CMB that chooses the arm randomly for each player. The bold scores are the best in each column.

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systemsccs: Computing methodologies Multi-agent reinforcement learning22footnotetext: $\dagger$ Corresponding author.
@@ 1. Introduction
Recommendation systems (RS) are widely deployed in various online platforms, such as Google, Facebook, and Yahoo!, to mitigate information overload. However, most existing recommendation methods (Rendle et al., 2009; He et al., 2020; Koren et al., 2009; Cheng et al., 2016; Zhang et al., 2021) only prioritize recommending the most relevant items to users, which can have negative consequences for both users and service providers. Users may experience the “filter bubble” (Pariser, 2011) problem, receiving recommendations based solely on their past behavior, leading to a lack of content diversity and reduced long-term satisfaction. Content providers may face the “Matthew Effect” (Merton, 1968), where new or niche content/products yield less exposure, discouraging the engagement of content providers. Therefore, it is crucial to consider not only relevance but also recommendation diversity to enhance the overall user experience and maintain a healthy ecosystem for content providers.
@@ 2.1. Recommendation Diversification
To address the “filter bubble” and reduced provider engagement issues, it is crucial to recommend accurate and diverse items for a healthier online marketplace. Existing diversification methods are typically offline and categorized as pre-processing, in-processing, and post-processing methods (Wu et al., 2024). Pre-processing methods (Kwon et al., 2020; Cheng et al., 2017; Paudel et al., 2017; Zheng et al., 2021) involve preparing interaction data before model training. In-processing integrates diversity into the training process, using it as regularization (Chen et al., 2020; Maropaki et al., 2020; Sanz-Cruzado and Castells, 2018; Wasilewski and Hurley, 2016; Zhou et al., 2020) or a ranking score (Li et al., 2017; Yu, 2022). Post-processing, the most scalable, includes greedy-based methods like MMR (Carbonell and Goldstein, 1998; Santos et al., 2010; Vargas et al., 2011) and DPP (Gan et al., 2020; Chen et al., 2018; Gong et al., 2022; Huang et al., 2021; Wilhelm et al., 2018), which adjust item selection and rankings to balance relevance and diversity, and refinement-based methods (Li et al., 2020b; Tsukuda and Goto, 2019), which modify positions or replace items based on diversity metrics. Other online methods, such as bandit strategies (Ding et al., 2021; Li et al., 2020a; Qin et al., 2014; Parapar and Radlinski, 2021), treat diversity as part of the score on each arm (item or topic) in the bandit recommendation algorithms and reinforcement learning (Stamenkovic et al., 2022; Zheng et al., 2018; Shi et al., 2023), continuously update based on user feedback for long-term optimization.
@@ 2.2. Explainable Recommendation
Explainable recommendations have attracted significant attention in academia and industry, aiming to enhance transparency, user satisfaction, and trust (Zhang et al., 2020, 2014, 2025). Early methods focused on generating individualized explanations, often customizing models and using auxiliary information (Zhang et al., 2014; Lu et al., 2018; Wu et al., 2019). For example, the Explicit Factor Model (EFM) (Zhang et al., 2014) recommends products based on features extracted from user reviews. Other approaches decouple explanations from the recommendation model, making them post-hoc and model-agnostic (Ribeiro et al., 2016; Wang et al., 2018; Singh and Anand, 2019; Chen et al., 2022). Recently, counterfactual reasoning has been widely used to improve explainability (Wachter et al., 2017; Ghazimatin et al., 2020; Tan et al., 2021; Xiong et al., 2021; Ge et al., 2022). For instance, PRINCE (Ghazimatin et al., 2020) identifies minimal user actions that, if removed, would change the recommendation, while CEF (Ge et al., 2022) uses counterfactual reasoning to explain fairness in feature-aware recommendation systems.
@@ 3.2. Counterfactual Framework for Explainable Diversification
Figure 1. The architecture of CMB. CMB consists of three major stages: the first stage of base model training, the second stage of counterfactual framework within multi-player bandits optimization, and the third stage of explanation. The $\mathbf{Q}$ is the learned item latent feature matrix from the base model $g$. The $\widetilde{\Psi}$ in the green part is alterable, which can be diversity or the trade-off between diversity and accuracy.
@@ 3.3. Multi-player Bandits for Diversity Optimization
When all players have selected the arm $\Delta_{p}^{t}$ , we can get the $\Delta$ and further the reward $r_{t}$ (Eq. 9) based on the counterfactual result ~$\widetilde{\mathcal{R}}_{g}$. To reduce the computational complexity, we update the cumulative reward value $V^{t+1}_{\Delta_{p}}$ of the corresponding arm $\Delta_{p}$ selected by each player $p$ in iteration $t+1$ by an incremental average method:
$V^{t+1}_{\Delta_{p}}$ $=\frac{1}{n}\sum_{i=1}^{n}r_{i},$
$=\frac{1}{n}\left(r_{n}^{t}+\sum_{i=1}^{n-1}r_{i}\right),$
$=\frac{1}{n}\left(r_{t}+\left(n-1\right)\frac{1}{n-1}\sum_{i=1}^{n-1}r_{i}\right),$
$=\frac{1}{n}\left(r_{t}+\left(n-1\right)V^{t}_{\Delta_{p}}\right),$
$=\frac{1}{n}\left(r_{t}+nV^{t}_{\Delta_{p}}-V^{t}_{\Delta_{p}}\right),$
(13) $=V^{t}_{\Delta_{p}}+\frac{1}{n}\left(r_{t}-V^{t}_{\Delta_{p}}\right),$
@@
Discussion. (1) We mainly work on the latent features of items to explain the recommendation diversification. The key consideration is that current mainstream recommendation models still originate from collaborative filtering methods, which are based on latent features. These motivate us to work on the latent features for controlling the diversity level of recommendation results. It is worth noting that our model can both work raw and latent features. (2) The main reason why we decided to apply the perturbation $\Delta$ to the features of items is that perturbing at the items’ feature level enables us to identify specific features that impact the diversity of the model at a more fine-grained feature level. (3) The multi-armed bandits method, unlike previous approaches, is primarily utilized in online recommendation scenarios. In our study, we employ this method to optimize the learning objective in the counterfactual framework, especially targeting and optimizing the non-differentiable diversity metrics directly.
@@ 4. Experiments
In this section, we conducted experiments on three real-world datasets to demonstrate the effectiveness of our method. We mainly focus on the following questions:
  • RQ1: Does our method get better recommendation diversification effects than the state-of-the-art methods? Especially the trade-off between recommendation accuracy and diversity.
  • RQ2: Do the selected top features play a significant role in diversification performance?
  • RQ3: Are our generated feature-level diversification explanations reasonable and intuitive in real cases?
  • RQ4: Do the hyperparameters ($A$, $n_{A}$, etc.) play a significant role in diversification performance?
@@ 4.1.1. Datasets
We performed experiments on three widely used real-world datasets, MovieLens 1M (Harper and Konstan, 2015) (ML1M), MovieLens 10M (Harper and Konstan, 2015) (ML10M), and Amazon CDs and Vinyl (Ni et al., 2019) (CDs) to evaluate the models under different data scales, data sparsity, and application scenarios. The statistics of the datasets are shown in Table 2.
@@ 4.1.4. Implementation Details
In our experiments, we employ BPRMF and LightGCN as the base model $g$, with LightGCN set to three layers of the graph neural network. For baselines, BPRMF calculates the relevance scores for MMR and constructs the kernel matrix for DPP. The trade-off parameter for MMR and DPP is empirically set to 0.9 after testing values from {0.1, 0.3, 0.5, 0.7, 0.9}, prioritizing guaranteeing the optimal accuracy performance as much as possible. All models have a latent feature dimension of 50, with $K$ set to 20 for counterfactual learning. The $\alpha$ in $\alpha$-nDCG is set to 0.5 same as (Santos et al., 2010; Vargas et al., 2011; Li et al., 2017; Qin et al., 2020; Yu, 2022), and model parameters are optimized by Adam (Kingma and Ba, 2014) with a learning rate of 0.005. In the bandit algorithm, we set the threshold of arm values $A$ to 0.3, the number of arms $n_{A}$ to 61, and $\epsilon$ to 0.1. The hyperparameters $\lambda_{1}$, $\lambda_{2}$, and the total number of iterations $T$ are set to 5, 0.9, and 200, respectively. The source code is available at https://github.com/Forrest-Stone/CMB.
@@
Table 3. Comparisons of the accuracy and diversity performance. The base model $g$ here adopts BPRMF. The bold scores are the best in each column, and the underlined scores are the second best. The symbols $\uparrow$ and $\downarrow$, along with their preceding values, represent the percentage (is omitted) improvement and decrease of a given method in the corresponding metric, in comparison to the base model $g$.
@@ 4.2. Performance Comparison (RQ1)
Second, not only does the combined optimization objective help achieve a reasonable balance between accuracy and diversity, but also the single diversity objective does. For instance, as shown in Table 5, when compared to BPRMF, $\text{CMB}_{\text{BPRMF}}$-ILAD shows a decrease of 18.79% in Recall@10 and an increase of 40.62% in ILAD@10 on ML10M, while showing a decrease of 31.29% in NDCG@10 and an increase of 59.58% in PC@10 on CDs. These results demonstrate that the proposed single diversity objective can improve diversity performance while maintaining accuracy performance, and outperforming other diversity methods like MMR and DPP.
@@
Table 4. Comparisons of the accuracy and diversity performance. The base model $g$ here adopts LightGCN. The bold scores are the best in each column, and the underlined scores are the second best. The symbols $\uparrow$ and $\downarrow$, along with their preceding values, represent the percentage (is omitted) improvement and decrease of a given method in the corresponding metric, in comparison to the base model $g$.
@@
Table 5. Comparisons among the accuracy and diversity performance of CMB optimizes the single diversity metrics. The base model $g$ here adopts BPRMF. $\text{CMB}_{\text{ILAD}}^{\text{Gradient}}$ represents CMB that directly optimizes the differentiable metric ILAD by using the gradient method. $\text{CMB}_{\text{BPRMF}}$-Random represents CMB that chooses the arm randomly for each player. The bold scores are the best in each column.
@@
Second, as illustrated in Tables 3 and 5, $\text{CMB}_{\text{BPRMF}}$-SC achieves the second-best SC@10 on ML10M, only 0.12% decrease compared to the best baseline. Similarly, $\text{CMB}_{\text{BPRMF}}$-PC also achieves the second-best performance in the PC metric on ML10M and CDs. These results demonstrate that CMB can achieve a satisfactory diversification. Moreover, while CMB sometimes trails DPP in diversity metrics, this is because DPP prioritizes diversity over accuracy. In contrast, our method balances both, leading to improved accuracy even if diversity scores are slightly lower when compared to the DPP.
Application observations. First, from Table 5, by comparing $\text{CMB}_{\text{BPRMF}}^{\text{Gradient}}$-ILAD and $\text{CMB}_{\text{BPRMF}}$-ILAD, we observe that the gradient descent optimization method outperforms the bandit optimization method in terms of accuracy on CDs, while exhibiting better diversity results on ML10M for PC and ILAD metrics. Thus, the choice of which optimization method should be based on the specific application scenario (e.g., dataset, diversity level, etc.).
Second, as shown in Table 3 and 5, compared to $\text{CMB}_{\text{Random}}$, CMB that only optimizes a single diversity metric also achieves good results on that metric, which shows that our bandit method is effective for optimizing different diversity metrics. Moreover, optimizing the combination of trade-off objectives simultaneously also achieves a good balance in accuracy and diversity, further highlighting the effectiveness of our combined optimization objectives. For example, NDCG@10 of $\text{CMB}_{\text{BPRMF}}$-SC-NDCG on ML10M and CDs is 18.88% and 34.08% higher than $\text{CMB}_{\text{BPRMF}}$-Random, respectively, while the SC@10 only decreases by 1.35% and 6.16%.
Other observations. First, the accuracy and diversity trade-off exists widely. No method can achieve the best results in both accuracy and diversity since an increase in accuracy generally corresponds to a decrease in diversity. From the results in Table 3, DPP achieves the best results in SC@10, PC@10, and ILAD@10 on ML1M and ML10M, but it achieves the worst performance in Recall@10 and NDCG@10 on these datasets.
Second, generally, no single method demonstrates superior performance across all diversity metrics. For example, as shown in Table 3, just consider diversity metrics, DPP has the highest SC@10, PC@10, and ILAD@10 performance on ML10M and CDs, but has the lowest $\alpha$-nDCG@K performance. This indicates the inherent gap between different diversity evaluation metrics, proving the necessity of optimizing different metrics in a general framework, which is just the focus of our work.
@@
4.3. Validity Analysis of Explanations (RQ2)
@@
4.4. Case Study of Explanations (RQ3)
@@ 5. Conclusion and Future Work
Acknowledgements.
This work was supported by the Early Career Scheme (No. CityU 21219323) and the General Research Fund (No. CityU 11220324) of the University Grants Committee (UGC), and the NSFC Young Scientists Fund (No. 9240127).
@@ A.2. Performance Comparison
Table 7. Comparisons of the accuracy and diversity performance. The base model $g$ here adopts BPRMF. The bold scores are the best in each column, and the underlined scores are the second best. The symbols $\uparrow$ and $\downarrow$, along with their preceding values, represent the percentage (is omitted) improvement and decrease of a given method in the corresponding metric, in comparison to the base model $g$.
@@
Table 8. Comparisons of the accuracy and diversity performance. The base model $g$ here adopts LightGCN. The bold scores are the best in each column, and the underlined scores are the second best. The symbols $\uparrow$ and $\downarrow$, along with their preceding values, represent the percentage (is omitted) improvement and decrease of a given method in the corresponding metric, in comparison to the base model $g$.
@@
Table 9. Comparisons among the accuracy and diversity performance of CMB optimizes the single diversity metrics. The base model $g$ here adopts BPRMF. $\text{CMB}_{\text{ILAD}}^{\text{Gradient}}$ represents CMB that directly optimizes the differentiable metric ILAD by using the gradient method. $\text{CMB}_{\text{BPRMF}}$-Random represents CMB that chooses the arm randomly for each player. The bold scores are the best in each column.